Solutions of sodium carbonate and silver nitrate react to form solid silver carbonate and a solution of sodium nitrate. A solution containing 3.20 g

g of sodium carbonate is mixed with one containing 4.57 g

g of silver nitrate. How many grams of sodium carbonate are present after the reaction is complete?


1
Expert's answer
2021-10-12T02:06:01-0400

Molar mass of sodium carbonate = 105.9888 g/mol

3.2/105.9888= 0.0302mol

Molar mass of silver nitrate= 169.87 g/mol

4.57/169.87= 0.027mol

0.027+0.0302= 0.0572mol

0.0572×105.9888= 6.063g